OpenStudy (anonymous):

how to find the eigenvector of this 1 0 1 0 1 0 1 0 1 I have found the eigenvalues which are 0, 1, 2

OpenStudy (amistre64):

lol, um, so how did you find them?

OpenStudy (amistre64):

-L along the diagonals; and take the determinant; set it equal to 0 to find Ls

OpenStudy (amistre64):

hard to tell what the vectors are without a vector to go by tho ...

OpenStudy (amistre64):

Ax = Lx Ax-Lx = 0 (A-L)x = 0 and solve for x

OpenStudy (amistre64):

subtract the Ls from the diags and rref to find "x"

OpenStudy (amistre64):

1-0 0 1 0 1-0 0 1 0 1-0 1-1 0 1 0 1-1 0 1 0 1-1 1-2 0 1 0 1-2 0 1 0 1-2

OpenStudy (amistre64):

rref{{1-0, 0, 1},{ 0, 1-0, 0},{ 1, 0, 1-0}} 1 0 1 x1 -1 0 1 0 x2 = x3 0 0 0 0 x3 1 rref{{1-1, 0, 1},{ 0, 1-1, 0},{ 1, 0, 1-1}} 100 0 001 = x3 0 000 1 rref{{1-2, 0, 1},{ 0, 1-2, 0},{ 1, 0, 1-2}} 10-1 1 01 0 = x3 0 00 0 1

OpenStudy (amistre64):

im not to sure about the second one tho

OpenStudy (anonymous):

I used the rule of sarus and got x^3-3x^2+2x (used x for lambda) and then i got 0, 1, 2 as eigenvalues

OpenStudy (amistre64):

ill trust you did that part good then ;)

OpenStudy (amistre64):

the rest is just row reducing that eugened matrixes with the 0 vector

OpenStudy (anonymous):

ok thanks :)